# Partially Retargeted Balancing Weights for Causal Effect Estimation Under Positivity Violations

### v1

@@ Discussion
$+\text{Var}(Z_{i}|\bm{X}_{i})E[\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}]^{2}$
$+\text{Var}[\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}]E[Z_{i}|\bm{X}_{i}]^{2}$
$=e(\bm{X}_{i})\{1-e(\bm{X}_{i})\}\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}+e(\bm{X}_{i})^{2}\{1-e(\bm{X}_{i})\}\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}$
@@
Then, we have that the asymptotic variance of the estimator when the propensity score model is misspecified is,
$E[\tilde{w}_{1}(\bm{X}_{i})^{2}e(\bm{X}_{i})\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}+\tilde{w}_{0}(\bm{X}_{i})^{2}\{1-e(\bm{X}_{i})\}\text{Var}\{Y_{i}(0)|\bm{X}_{i}\}+\{\tau(\bm{X}_{i})-\tau\}^{2}]$
@@
We explore the restrictions of the implied propensity score model commonly used measures of dispersion: 1) entropy weights; and 2) stable balancing weights. The first order conditions in (7) imply the following condition:
@@
Without a defining a specific measure of dispersion $D(\cdot)$ (and therefore a specific $\rho(\cdot)$) it is hard to connect this to $Var(Z|\bm{X})$. However, we can explore two common deviance measures (Wang and Zubizarreta,, 2020):
@@
    Entropy: $D(t)=t\log(t)$ $\implies$ $\rho^{\prime}(t)=e^{-t-1}$ $\implies$ $(\rho^{\prime})^{-1}(t)=-\log(t)-1$
@@
Consider the entropy measure of weight dispersion first. Applying this to the above equations, we get the following
@@
such that $\text{Var}(Z|\bm{X}=\bm{x})$ is modeled as only varying with $c(\bm{x})$. Now, consider the dispersion measure used in the stable balancing weights paper:
@@
and again, $\text{Var}(Z|\bm{X}=\bm{x})$ is modeled as only varying with $c(\bm{x})$. While this relationship may not hold for all $D(\cdot)$ functions, we can generally think of this formulation of the direct balancing problem as allowing us to model $E[Z|\bm{X}=\bm{x}]$ with all functions of covariates, but $\text{Var}(Z|\bm{X}=\bm{x})$ with only the set of covariates that are balanced to the population. However, since this in general does not hold for an arbitrary function $g(\bm{x})$, this places substantial restrictions on what forms $g(\bm{x})$ can take for the implied propensity score model to be true. This $g(\bm{x})$ form can be identified for each $D(\cdot)$ by using the first order conditions (7a)-(7c) to solve for $\bm{\gamma}^{T}g(\bm{x})$ as a function of $\bm{\alpha}_{z}^{T}c(\bm{x})$.
@@
Consider the case where there are $M>2$ treatment groups such that $Z_{i}=m$ for $m=0,\ldots,M-1$ and where we want to estimate the average treatment effect between treatments $m_{1},m_{2}=0,\ldots,M-1$, $\tau_{m_{1},m_{2}}=E[\mu_{m_{1}}(\bm{x})-\mu_{m_{2}}(\bm{x})]$. Adapting Problem (5) to derive weights for estimating $\tau_{m_{1},m_{2}}$ simply requires replacing $z=0,1$ with $z=m_{1},m_{2}$ in Problem (5). However, if weights are derived in this way for a series of average treatment effects (e.g., $\tau_{m_{1},m_{2}},\tau_{m_{2},m_{3}},\tau_{m_{1},m_{3}}$ etc.,), each average treatment effect will correspond to a different population for the covariate set $g(\bm{x})$ or even a different set of covariates in $g(\bm{x})$. To maintain the same $g(\bm{x})$ covariate population for average treatment effects corresponding to a set all potential contrasts for $M_{1}\leq M$ treatments, we would include $M_{1}$ constraints balancing $c(\bm{x})$ in treatment groups to the sample population (e.g., constraint (5b)) and $M_{1}-1$ constraints balancing $g(\bm{x})$ between treatment groups (e.g., constraint (5c)). As $M_{1}$ increases, generally the set $g(\bm{x})$ will need to be larger for the existence of a solution. However, for there to be no estimator error, $g(\bm{x})$ must contain no effect modifiers for all treatment groups. Thus, deriving weights for all contrasts separately may allow for a smaller $g(\bm{x})$ set, but this set and corresponding population may vary for each contrast. In contrast, deriving weights for contrasts together will result in the same target population for all average treatment effect estimands but the set $g(\bm{x})$ may need to be large to derive the weights.
@@
Consider the estimand $E[h(\bm{X})\tau(\bm{X})]$ for some function $h(\bm{x})$. This as a weighted average treatment effect (WATE) where the population covariate density $f(\bm{x})$ is modified to $h(\bm{x})f(\bm{x})$. The ATO is the WATE when $h(\bm{x})=e(\bm{x})\{1-e(\bm{x})\}$. When adapting our proposed constrained optimization problem to derive weights for estimating the WATE, $\{w_{i}^{WATE}\}$, we can simply replace the right-hand side of constraint (5b) with $\sum_{i=1}^{n}h(\bm{X}_{i})c_{k}(\bm{X}_{i})/\sum_{i=1}^{n}h(\bm{X}_{i})$. The resulting weights will balance $c(\bm{x})$ to the weighted population with density $h(\bm{x})f(\bm{x})$ while $g(\bm{x})$ will only be balanced between the treated and control group. Similar to the ATT, all discussion and methods proposed in Sections 3.2 and 6 can be used for the WATE. However, if the population defined by $h(\bm{x})f(\bm{x})$ has improved overlap compared to the sample population, the set $g(\bm{x})$ required to yield a solution will be smaller than the set required to yield a solution for the ATE. Thus, various $h(\bm{x})$ functions could be used in combination with our proposed constrained optimization problem to improve overlap and yield a solution. Yet, this will modify a large portion of the sample covariate population to the population described by $h(\bm{x})f(\bm{x})$ such that this modified population ideally is of specific interest to the given research question.

### v2

@@ Discussion
$\quad+\text{Var}(Z_{i}|\bm{X}_{i})E[\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}]^{2}$
$\quad+\text{Var}[\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}|\bm{X}_{i}]E[Z_{i}|\bm{X}_{i}]^{2}$
$=e(\bm{X}_{i})\{1-e(\bm{X}_{i})\}\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}$
$\quad+e(\bm{X}_{i})^{2}\{1-e(\bm{X}_{i})\}\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}$
@@
Then, we have that the asymptotic variance of the estimator when the propensity score model is misspecified is
$E[\tilde{w}_{1}(\bm{X}_{i})^{2}e(\bm{X}_{i})\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}+\tilde{w}_{0}(\bm{X}_{i})^{2}\{1-e(\bm{X}_{i})\}\text{Var}\{Y_{i}(0)|\bm{X}_{i}\}+\{\tau(\bm{X}_{i})-\tau\}^{2}].$
Corollary 4.2
Proof.
Assume that $\mu_{z}(\bm{x})$ and $\tau(\bm{x})$ are linear in both $c(\bm{x})$ and $g(\bm{x})$ such that $\mu_{1}(\bm{x})=\bm{\beta}_{1}^{T}c(\bm{x})+\bm{\lambda}_{1}^{T}g(\bm{x})$ and $\mu_{0}(\bm{x})=\bm{\beta}_{0}^{T}c(\bm{x})+\bm{\lambda}_{0}^{T}g(\bm{x})$ for some $\bm{\beta}_{z}\in\mathbb{R}^{K}$ and $\bm{\lambda}_{z}\in\mathbb{R}^{L}$ where $\bm{\beta}_{0}\neq\bm{\beta}_{1}$ and $\bm{\lambda}_{0}\neq\bm{\lambda}_{1}$. Instead of targeting the ATE, $\tau$, $\hat{\tau}_{w^{*}}$ targets $\tau_{g,\tilde{w}}=E[E[Y|Z=1,c(\bm{X}),\tilde{w}(\bm{X})e(\bm{X})g(\bm{X_{i}})]]-E[E[Y|Z=0,c(\bm{X}),\tilde{w}(\bm{X})\{1-e(\bm{X})\}g(\bm{X})]]=E[\tau_{g,\tilde{w}}(\bm{X},Z)]$ where $\tau_{g,\tilde{w}}(\bm{x},z)=(\bm{\beta}_{1}-\bm{\beta}_{0})^{T}c(\bm{x})+\bm{\lambda}_{1}^{T}\tilde{w}_{1}(\bm{x})zg(\bm{x})-\bm{\lambda}_{0}^{T}\tilde{w}_{0}(\bm{x})(1-z)g(\bm{x})$. We have the following decomposition of estimator error:
$\hat{\tau}_{w^{*}}-\tau_{g,\tilde{w}}$ $=\frac{1}{n}\sum_{i=1}^{n}w^{*}_{1}(\bm{X}_{i})Z_{i}Y_{i}-\frac{1}{n}\sum_{i=1}^{n}w^{*}_{0}(\bm{X}_{i})(1-Z_{i})Y_{i}-\tau_{g,\tilde{w}}$
$=\frac{1}{n}\sum_{i=1}^{n}\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}-\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}-\tau_{g,\tilde{w}}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}\{w_{1}^{*}(\bm{X}_{i})-\tilde{w}_{1}(\bm{X}_{i})\}Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}$
$\quad-\frac{1}{n}\sum_{i=1}^{n}\{w_{0}^{*}(\bm{X}_{i})-\tilde{w}_{0}(\bm{X}_{i})\}(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}w^{*}_{1}(\bm{X}_{i})Z_{i}\mu_{1}(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}w^{*}_{0}(\bm{X}_{i})(1-Z_{i})\mu_{0}(\bm{X}_{i})$
$=\frac{1}{n}\sum_{i=1}^{n}\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}-\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}\bm{\beta}_{1}^{T}c(\bm{X}_{i})-\bm{\beta}_{0}^{T}c(\bm{X}_{i})$
$\quad\quad+E[\bm{\lambda}_{1}^{T}\tilde{w}_{1}(\bm{X})Zg(\bm{X})]-E[\bm{\lambda}_{0}^{T}\tilde{w}_{0}(\bm{X})(1-Z)g(\bm{X})]-\tau_{g,\tilde{w}}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}\{w_{1}^{*}(\bm{X}_{i})-\tilde{w}_{1}(\bm{X}_{i})\}Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}$
$\quad-\frac{1}{n}\sum_{i=1}^{n}\{w_{0}^{*}(\bm{X}_{i})-\tilde{w}_{0}(\bm{X}_{i})\}(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}\bm{\lambda}_{1}^{T}w^{*}_{1}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})-E[\bm{\lambda}_{1}^{T}\tilde{w}_{1}(\bm{X})Zg(\bm{X})]$
$\quad-\frac{1}{n}\sum_{i=1}^{n}\bm{\lambda}_{0}^{T}w^{*}_{0}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})+E[\bm{\lambda}_{0}^{T}\tilde{w}_{0}(\bm{X})(1-Z)g(\bm{X})].$
By Assumption Assumption 3: Balance function conditions and the assumption that $\rho^{\prime\prime}\{\cdot\}$ and $\rho^{\prime\prime\prime}\{\cdot\}$ are continuous over $\mathbb{R}$, the results of Theorems 5.28 and 7.2 in Boos and Stefanski (2013) apply. Thus, there exist some $q_{1}(\bm{X}_{i},Z_{i})$ and $q_{0}(\bm{X}_{i},Z_{i})$ with $E[q_{z}(\bm{X}_{i},Z_{i})]=0$, $z=0,1$ such that
$\frac{1}{n}\sum_{i=1}^{n}\bm{\lambda}_{1}^{T}w^{*}_{1}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})-E[\bm{\lambda}_{1}^{T}\tilde{w}_{1}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})]$ $=\frac{1}{n}\sum_{i=1}^{n}q_{1}(\bm{X}_{i},Z_{i})+R_{2}$
$\frac{1}{n}\sum_{i=1}^{n}\bm{\lambda}_{0}^{T}w^{*}_{0}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})-E[\bm{\lambda}_{0}^{T}\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})]$ $=\frac{1}{n}\sum_{i=1}^{n}q_{0}(\bm{X}_{i},Z_{i})+R_{3}$
where $R_{2},R_{3}=o_{p}(n^{-1/2})$. Note that we know the exact form of these two functions, $q_{1}(\bm{X}_{i},Z_{i})$ and $q_{0}(\bm{X}_{i},Z_{i})$, as specified in Theorem 5.28 in Boos and Stefanski (2013). Then $\hat{\tau}_{w}-\tau_{g,\tilde{w}}=S_{i}+R_{1}+R_{2}+R_{3}$ where
$S_{i}$ $=\tilde{w}_{1}(\bm{X}_{i})Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}-\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}$
$\quad+\bm{\beta}_{1}^{T}c(\bm{X}_{i})-\bm{\beta}_{0}^{T}c(\bm{X}_{i})-E[\bm{\beta}_{1}^{T}c(\bm{X}_{i})-\bm{\beta}_{0}^{T}c(\bm{X}_{i})]+q_{1}(\bm{X}_{i},Z_{i})-q_{0}(\bm{X}_{i},Z_{i})$
$R_{1}$ $=\frac{1}{n}\sum_{i=1}^{n}\{w_{1}^{*}(\bm{X}_{i})-\tilde{w}_{1}(\bm{X}_{i})\}Z_{i}\{Y_{i}(1)-\mu_{1}(\bm{X}_{i})\}$
$\quad-\frac{1}{n}\sum_{i=1}^{n}\{w_{0}^{*}(\bm{X}_{i})-\tilde{w}_{0}(\bm{X}_{i})\}(1-Z_{i})\{Y_{i}(0)-\mu_{0}(\bm{X}_{i})\}$
where $R_{1}$ has previously been shown to be $o_{p}(n^{-1/2})$. Then,
[…]
$=E[\tilde{w}_{1}(\bm{X}_{i})^{2}e(\bm{X}_{i})\text{Var}\{Y_{i}(1)|\bm{X}_{i}\}+\tilde{w}_{0}(\bm{X}_{i})^{2}\{1-e(\bm{X}_{i})\}\text{Var}\{Y_{i}(0)|\bm{X}_{i}\}]$
[…]
$\quad\quad\quad\quad\left.\left.+q_{1}(\bm{X}_{i},Z_{i})-q_{0}(\bm{X}_{i},Z_{i})\right\}^{2}\right]$
as all covariance terms cancel as shown in Corollary 4.1. It remains to determine $q_{1}(\bm{X}_{i},Z_{i})$ and $q_{0}(\bm{X}_{i},Z_{i}).$ By Theorem 5.28 in Boos and Stefanski (2013),
$q_{1}(\bm{X}_{i},Z_{i})$ $=\bm{\lambda}_{1}^{T}\tilde{w}_{1}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})-E[\bm{\lambda}_{1}^{T}\tilde{w}_{1}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})]+r_{1}(\bm{X}_{i},Z_{i})\>\>\>\text{and}$
$q_{0}(\bm{X}_{i},Z_{i})$ $=\bm{\lambda}_{0}^{T}\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})-E[\bm{\lambda}_{0}^{T}\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})]+r_{0}(\bm{X}_{i},Z_{i}),$
where $r_{z}(\bm{X}_{i},Z_{i})$ depend on the asymptotic distribution of $w^{*}_{1}(\bm{X}_{i})$ and $w^{*}_{0}(\bm{X}_{i})$. Therefore, to derive $r_{z}(\bm{X}_{i},Z_{i})$, we first derive the asymptotic distribution of $\hat{\bm{\alpha}}_{1},\hat{\bm{\alpha}}_{0},\hat{\bm{\gamma}}$ using the theory of M-estimation.
Let $m_{k}=E[c_{k}(\bm{X}_{i})]$ and $\bm{m}=(m_{1},\ldots m_{K})$. Let $\bm{\theta}=(\bm{m},\bm{\alpha}_{1},\bm{\alpha}_{0},\bm{\gamma})$. Let
$\phi_{k}(\bm{X},\bm{m})$ $=c_{k}(\bm{X})-m_{k}$
$\varphi^{(1)}_{k}(\bm{X},Z,\bm{m},\bm{\alpha}_{1},\bm{\gamma})$ $=\rho^{\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X})+\bm{\gamma}^{T}g(\bm{X})\}Zc_{k}(\bm{X})-m_{k}$
$\varphi^{(0)}_{k}(\bm{X},Z,\bm{m},\bm{\alpha}_{0},\bm{\gamma})$ $=\rho^{\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X})-\bm{\gamma}^{T}g(\bm{x})\}(1-Z)c_{k}(\bm{X})-m_{k}$
$\varphi^{(2)}_{l}(\bm{X},Z,\bm{\alpha}_{0},\bm{\alpha}_{1}\bm{\gamma})$ $=\rho^{\prime}\{\bm{\alpha}_{1}^{T}c(\bm{X})+\bm{\gamma}^{T}g(\bm{X})\}Zg_{l}(\bm{X})-\rho^{\prime}\{\bm{\alpha}_{0}^{T}c(\bm{X})-\bm{\gamma}^{T}g(\bm{X})\}(1-Z)g_{l}(\bm{X})$
Let $\psi(\bm{X},Z,\bm{\theta})=(\phi,\varphi^{(1)},\varphi^{(0)},\varphi^{(2)})$. Then $\hat{\alpha}_{0},\hat{\alpha}_{1},\hat{\gamma}$ is the solution to
$\frac{1}{n}\sum_{i=1}^{n}\psi(\bm{X}_{i},Z_{i},\bm{\theta})=0.$
Given the assumptions on $\rho\{\cdot\}$, Assumption Assumption 3: Balance function conditions, and similar arguments to the proof of Lemma 6, by the theory of M-estimation and Theorem 7.2 in Boos and Stefanski (2013),
$\hat{\bm{\theta}}-\tilde{\bm{\theta}}$ $=\sum_{i=1}^{n}\bm{A}^{-1}(\tilde{\bm{\theta}})\psi(\bm{X}_{i},Z_{i},\bm{\tilde{\theta})}+R_{4}$
[…]
$=\begin{bmatrix}\bm{I}_{K}&\bm{0}_{K}&\bm{0}_{K}&\bm{0}_{K\times L}\\ \bm{I}_{K}&-\bm{H}_{h^{\prime}_{1}c(\bm{X}),c(\bm{X})}&\bm{0}_{K}&-\bm{H}_{h^{\prime}_{1}c(\bm{X}),g(\bm{X})}\\ \bm{I}_{K}&\bm{0}_{K}&-\bm{H}_{h^{\prime}_{0}c(\bm{X}),c(\bm{X})}&-\bm{H}_{h^{\prime}_{0}c(\bm{X}),g(\bm{X})}\\ \bm{0}_{L\times K}&-\bm{H}_{h^{\prime}_{1}g(\bm{X}),c(\bm{X})}&\bm{H}_{h^{\prime}_{0}g(\bm{X}),c(\bm{X})}&-\bm{H}_{(h^{\prime}_{1}+h^{\prime}_{0})g(\bm{X}),g(\bm{X})}\\ \end{bmatrix}$
[…]
$q_{1}(\bm{X}_{i},Z_{i})$ $=\bm{\lambda}_{1}^{T}\tilde{w}_{1}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})-E[\bm{\lambda}_{1}^{T}\tilde{w}_{1}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})]$
$\quad+E\left[\frac{\partial}{\partial(\tilde{\bm{\alpha}}_{1},\tilde{\bm{\gamma}})}\bm{\lambda}_{1}^{T}\tilde{w}_{1}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})\right]\bm{A}^{-1}(\tilde{\bm{\theta}})\psi(\bm{X}_{i},Z_{i},\bm{\tilde{\theta})}_{(\tilde{\bm{\alpha}}_{1},\tilde{\bm{\gamma}})}$
$q_{0}(\bm{X}_{i},Z_{i})$ $=\bm{\lambda}_{0}^{T}\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})-E[\bm{\lambda}_{0}^{T}\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})]$
$\quad+E\left[\frac{\partial}{\partial(\tilde{\bm{\alpha}}_{0},\tilde{\bm{\gamma}})}\bm{\lambda}_{0}^{T}\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})\right]\bm{A}^{-1}(\tilde{\bm{\theta}})\psi(\bm{X}_{i},Z_{i},\bm{\tilde{\theta})}_{(\tilde{\bm{\alpha}}_{0},\tilde{\bm{\gamma}})}$,
where
$E\left[\frac{\partial}{\partial(\tilde{\bm{\alpha}}_{1},\tilde{\bm{\gamma}})}\bm{\lambda}_{1}^{T}\tilde{w}_{1}(\bm{X}_{i})Z_{i}g(\bm{X}_{i})\right]$ $=[\bm{\lambda}_{1}^{T}\bm{H}_{h^{\prime}_{1}g(\bm{X}),c(\bm{X})},\bm{\lambda}_{1}^{T}\bm{H}_{h^{\prime}_{1}g(\bm{X}),g(\bm{X})}]\text{ and}$
$E\left[\frac{\partial}{\partial(\tilde{\bm{\alpha}}_{0},\tilde{\bm{\gamma}})}\bm{\lambda}_{0}^{T}\tilde{w}_{0}(\bm{X}_{i})(1-Z_{i})g(\bm{X}_{i})\right]$ $=[\bm{\lambda}_{0}^{T}\bm{H}_{h^{\prime}_{0}g(\bm{X}),c(\bm{X})},-\bm{\lambda}_{0}^{T}\bm{H}_{h^{\prime}_{0}g(\bm{X}),g(\bm{X})}].$
Then, we have that
[…]
$\quad+E[\{\tau_{g,\tilde{w}}(\bm{X}_{i},Z_{i})-\tau_{g,\tilde{w}}+r_{1}(\bm{X}_{i},Z_{i})-r_{0}(\bm{X}_{i},Z_{i})\}^{2}]$
$=V_{g,\tilde{w}}$
for $r_{1}(\bm{X}_{i},Z_{i})=[\bm{\lambda}_{1}^{T}\bm{H}_{h^{\prime}_{1}g(\bm{X}),c(\bm{X})},\bm{\lambda}_{1}^{T}\bm{H}_{h^{\prime}_{1}g(\bm{X}),g(\bm{X})}]\bm{A}^{-1}(\tilde{\bm{\theta}})\psi(\bm{X}_{i},Z_{i},\bm{\tilde{\theta})}_{(\tilde{\bm{\alpha}}_{1},\tilde{\bm{\gamma}})}$ and $r_{0}(\bm{X}_{i},Z_{i})=[\bm{\lambda}_{0}^{T}\bm{H}_{h^{\prime}_{0}g(\bm{X}),c(\bm{X})},-\bm{\lambda}_{0}^{T}\bm{H}_{h^{\prime}_{0}g(\bm{X}),g(\bm{X})}]\bm{A}^{-1}(\tilde{\bm{\theta}})\psi(\bm{X}_{i},Z_{i},\bm{\tilde{\theta})}_{(\tilde{\bm{\alpha}}_{0},\tilde{\bm{\gamma}})}$. Thus,
[…]
as desired. ∎
@@
We explore the restrictions on the implied propensity score model for two commonly used measures of dispersion: 1) entropy weights (Hainmueller, 2012); and 2) stable balancing weights (Zubizarreta, 2015). The first order conditions in (7) imply the following condition:
@@
Without a defining a specific measure of dispersion $D(\cdot)$ (and therefore a specific $\rho(\cdot)$) it is hard to connect this to $\text{Var}(Z|\bm{X})$. However, we explore two common deviance measures (Wang and Zubizarreta, 2020):
@@
    Entropy weights: $D(t)=t\log(t)$ $\implies$ $\rho^{\prime}(t)=e^{-t-1}$ $\implies$ $(\rho^{\prime})^{-1}(t)=-\log(t)-1$
@@
Consider the entropy measure of weight dispersion first. Applying this to the first order conditions in (7), we get the following
@@
such that $\text{Var}(Z|\bm{X}=\bm{x})$ only varies by $c(\bm{x})$ in the implied propensity score model. Now, consider the dispersion measure used for stable balancing weights:
@@
and again, $\text{Var}(Z|\bm{X}=\bm{x})$ only varies by $c(\bm{x})$ in the implied propensity score model. While this relationship may not hold for every possible dispersion measure, for these two dispersion measures the implied propensity score model of Problem (5) allows $E[Z|\bm{X}=\bm{x}]$ to vary with all functions of covariates but $\text{Var}(Z|\bm{X}=\bm{x})$ to vary only with the set of covariates that are balanced to the population. Since this relationship does not hold for any arbitrary $c(\bm{x})$ and $g(\bm{x})$ sets, this model places substantial restrictions on what forms these sets can take for the implied propensity score model to be true. Given a specific $c(\bm{x})$ set, the $g(\bm{x})$ set that satisfies the implied propensity score model can be identified for each $D(\cdot)$ by using the first order conditions (7a)-(7c) to solve for $\bm{\gamma}^{T}g(\bm{x})$ as a function of $\bm{\alpha}_{z}^{T}c(\bm{x})$.
@@
Consider the case where there are $M>2$ treatment groups such that $Z_{i}=m$ for $m=0,\ldots,M-1$ and where we want to estimate the average treatment effect between treatments $m_{1},m_{2}=0,\ldots,M-1$, $\tau_{m_{1},m_{2}}=E[\mu_{m_{1}}(\bm{x})-\mu_{m_{2}}(\bm{x})]$. Adapting Problem (5) to derive weights for estimating $\tau_{m_{1},m_{2}}$ simply requires replacing $z=0,1$ with $z=m_{1},m_{2}$ in Problem (5). However, if weights are derived in this way for a series of average treatment effects (e.g., $\tau_{m_{1},m_{2}},\tau_{m_{2},m_{3}},\tau_{m_{1},m_{3}}$ etc.,), each average treatment effect will correspond to a different population for the covariate set $g(\bm{x})$ or even a different $g(\bm{x})$ set. To maintain the same $g(\bm{x})$ covariate population for average treatment effects corresponding all contrasts for $M_{1}\leq M$ treatments, we would include $M_{1}$ constraints balancing $c(\bm{x})$ in treatment groups to the sample population (e.g., constraint (5b)) and $M_{1}-1$ constraints balancing $g(\bm{x})$ between treatment groups (e.g., constraint (5c)). As $M_{1}$ increases, generally the set $g(\bm{x})$ will need to be larger for the existence of a solution. However, for there to be no estimator error for every contrast, $g(\bm{x})$ must contain no effect modifiers for all treatment groups. Thus, deriving weights for all contrasts separately may allow for a smaller $g(\bm{x})$ set, but this set and corresponding population may vary for each contrast. In contrast, deriving weights for all contrasts together will result in the same target population for all average treatment effect estimands but the set $g(\bm{x})$ may need to be large to derive the weights.
@@
Consider the estimand $E[h(\bm{X})\tau(\bm{X})]$ for some function $h(\bm{x})$. This as a weighted average treatment effect (WATE) where the population covariate density $f(\bm{x})$ is modified to $h(\bm{x})f(\bm{x})$. The ATO is the WATE when $h(\bm{x})=e(\bm{x})\{1-e(\bm{x})\}$. When adapting our proposed constrained optimization problem to derive weights for estimating the WATE, $\{w_{i}^{WATE}\}$, we can simply replace the right-hand side of constraint (5b) with $\sum_{i=1}^{n}h(\bm{X}_{i})c_{k}(\bm{X}_{i})/\sum_{i=1}^{n}h(\bm{X}_{i})$. The resulting weights will balance $c(\bm{x})$ to the weighted population with density $h(\bm{x})f(\bm{x})$ while $g(\bm{x})$ will only be balanced between the treated and control groups. Similar to the ATT, all discussion and methods proposed in Sections 3.2 and 5 can be used for the WATE. However, if the population defined by $h(\bm{x})f(\bm{x})$ has improved overlap compared to the sample population, the set $g(\bm{x})$ required to yield a solution will be smaller than the set required to yield a solution for the ATE. Thus, various $h(\bm{x})$ functions could be used in combination with our proposed constrained optimization problem to improve overlap and yield a solution. Yet, this will modify a large portion of the sample covariate population to the population described by $h(\bm{x})f(\bm{x})$ such that this modified population ideally is of specific interest to the given research question.